# Reward-free World Models for Online Imitation Learning

### v1

@@ 5.3 Ablation Studies
In this section, we will show the ablation studies of our model, including the ablation over expert trajectory numbers. Regarding the ablations for objective formulation and gradient penalty selection, we refer to Appendix E.3.
@@
6 Conclusions and Broader Impact
We propose an online imitation learning approach that utilizes reward-free world models to address tasks in complex environments. By incorporating latent planning and dynamics learning, our model can have a deeper understanding of intricate environment dynamics. We demonstrate stable, expert-level performance on challenging tasks, including dexterous hand manipulation and high-dimensional locomotion control. In terms of broader impact, our model holds potential for real-world applications in manipulation and locomotion, particularly for tasks that involve visual inputs and complex environment dynamics.
@@ E.1 Additional High-dimensional Locomotion Experiments
Figure 10: Additional High-dimensional Locomotion Experiments Our IQ-MPC shows stable and expert-level performance on different tasks in the Dog environment, which demonstrates our model’s capability in handling high-dimensional tasks. In the plots, the color settings are the same as those in Figure 2.
@@ E.3 Additional Ablation Studies
In this section, we demonstrate the results of ablating over objective formulation and gradient penalty.
@@
We observe performance improvement using the reformulated objective Eq.12 as we mentioned in the Model Learning part of Section 4.1. In details, we changed the value temporal difference term $\mathbb{E}_{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}^{\prime}_{t})\sim\mathcal{B}_{\pi}}[V^{\pi}(\mathbf{z}_{t})-\gamma V^{\pi}(\mathbf{z}^{\prime}_{t})]$ into a form only containing value from initial distribution $\mathbb{E}_{\mathbf{s}_{0}\sim\mathcal{B}_{E}}[(1-\gamma)V^{\pi}(\mathbf{z}_{0})]$. This technique is also mentioned in the original IQ-Learn paper (Garg et al., 2021). We have given the theoretical proof for mathematical equivalence in Theorem 1. In this section, we provide the empirical analysis regarding the effectiveness of this technique in the context of our IQ-MPC model. We observe stabler Q estimation leveraging this technique. Moreover, in this case, the difference in Q estimation between the expert batch and the behavioral batch can converge more easily, especially for high-dimensional cases like the Humanoid Walk and Dog Run task. The better convergence of Q estimation difference shows that the Q function faces difficulty in distinguishing between expert and behavioral demonstrations, which implies that the policy prior behaves similarly as expert demonstrations. The stable Q estimation results in a better learning behavior for the latent dynamics model, which is observed by measuring the prediction consistency loss (The first term in Eq.11) during training. The results in Humanoid Walk task are shown in Figure 12.
@@
Appendix F Additional Theoretical Supports
@@
F.1 Objective Equivalence
In this section, we will provide proof for the reformulation of the second term in Eq.12 for completeness. We borrow the proof from Garg et al. (2021) and slightly modify it to fit our setting with latent representations instead of actual states. The proof is demonstrated in Theorem 1. In Eq.12, we use the mean over encoded latent representation batch sampled from the expert buffer $\mathcal{B}_{E}$ to approximate the mean over initial distribution $\tilde{p}_{0}$ on latent representation.
Theorem 1 (Objective Equivalence).
@@ F.1 Objective Equivalence
F.2 Policy Update Guarantee

### v2

@@ 5.2 Ablation Studies
In this section, we will show the ablation studies of our model, including the ablation over expert trajectory numbers. Regarding the ablations for objective formulation, gradient penalty selection, and hyperparameter $\alpha$, we refer to Appendix E.3.
@@
6 Conclusions
We propose an online imitation learning approach using reward-free world models for complex tasks. With latent planning and dynamics learning, our model achieves expert-level performance in dexterous hand manipulation and high-dimensional locomotion. It has potential for real-world applications involving visual inputs and intricate environment dynamics.
@@ E.1 Additional High-dimensional Locomotion Experiments
Figure 8: Additional High-dimensional Locomotion Experiments Our IQ-MPC shows stable and expert-level performance on different tasks in the Dog environment, which demonstrates our model’s capability in handling high-dimensional tasks. In the plots, the blue lines and orange lines represent the results from IQL+SAC (Garg et al., 2021) and CFIL+SAC (Freund et al., 2023), respectively, while the red lines represent the results from our IQ-MPC.
@@ E.3 Additional Ablation Studies
In this section, we demonstrate the results of ablating over objective formulation, gradient penalty, and hyperparameter $\alpha$.
@@
We observe performance improvement using the reformulated objective Eq.12 as we mentioned in the Model Learning part of Section 4.1. In details, we changed the value temporal difference term $\mathbb{E}_{(\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}^{\prime}_{t})\sim\mathcal{B}_{\pi}}[V^{\pi}(\mathbf{z}_{t})-\gamma V^{\pi}(\mathbf{z}^{\prime}_{t})]$ into a form only containing value from initial distribution $\mathbb{E}_{\mathbf{s}_{0}\sim\mathcal{B}_{E}}[(1-\gamma)V^{\pi}(\mathbf{z}_{0})]$. This technique is also mentioned in the original IQ-Learn paper (Garg et al., 2021). We have given the theoretical proof for mathematical equivalence in Lemma 2. In this section, we provide the empirical analysis regarding the effectiveness of this technique in the context of our IQ-MPC model. We observe stabler Q estimation leveraging this technique. Moreover, in this case, the difference in Q estimation between the expert batch and the behavioral batch can converge more easily, especially for high-dimensional cases like the Humanoid Walk and Dog Run task. The better convergence of Q estimation difference shows that the Q function faces difficulty in distinguishing between expert and behavioral demonstrations, which implies that the policy prior behaves similarly as expert demonstrations. The stable Q estimation results in a better learning behavior for the latent dynamics model, which is observed by measuring the prediction consistency loss (The first term in Eq.11) during training. The results in Humanoid Walk task are shown in Figure 10.
@@
Hyperparameter Selection
We perform an ablation study on the selection of the hyperparameter $\alpha$ in Eq.12. The hyperparameter $\alpha$ controls the strength of the $\chi^{2}$ regularization applied to the inverse soft-Q objective. Intuitively, the last term in Eq.12 serves as a penalty on the magnitude of the estimated reward. Therefore, smaller values of $\alpha$ result in a larger penalty on the estimated reward magnitude, which helps enforce training stability and prevents Q estimation from exploding. In contrast, larger values of $\alpha$ encourage more aggressive estimation of the reward and Q value, increasing the chances of training instability. We experiment with the effect of this hyperparameter in the Humanoid Walk task and conclude that $\alpha=0.5$ is the optimal choice. We present our results in Figure 12.
Figure 12: Ablation on Hyperparameter Selection. In our ablation study, we found that larger values of $\alpha$ may lead to higher Q estimations, resulting in suboptimal and unstable training behavior. Conversely, smaller values of $\alpha$ lead to lower Q estimations, which correspond to stronger regularization and may also cause suboptimal performance.
E.4 Visual Experiments
E.5 Experiments on Noisy Environment Dynamics
In this section, we evaluate the robustness of our IQ-MPC model under noisy and stochastic environment dynamics. HyPE (Ren et al., 2024) has demonstrated relatively robust performance when subjected to minor noise perturbations in environment transitions. Although our model is primarily designed for fully deterministic settings, we observe that it exhibits a degree of robustness when handling stochastic environment dynamics.
For our experiments, we adopt the same environment settings as HyPE (Ren et al., 2024), introducing a trembling noise probability, $p_{tremble}$, into the environment transitions. Specifically, we assess the impact of $p_{tremble}$ on our IQ-MPC model in the Walker Run task. The results indicate that our model maintains certain robustness even in the presence of transition noise. We present the results the Figure 14.
Figure 14: Experiments on Noisy Environment Dynamics We evaluate our model’s performance on the Walker Run task under different values of $p_{tremble}$, where a larger $p_{tremble}$ indicates greater stochasticity in the environment dynamics. Specifically, we experiment with $p_{tremble}=0.01$ and $p_{tremble}=0.02$, observing only slight performance degradation. This suggests that our model exhibits a degree of robustness to noisy environment dynamics.
E.6 Reward Recovery
We further analyze the correlation between the decoded rewards and ground-truth rewards in the Hopper Hop, Cheetah Run, Quadruped Run, and Walker Run tasks. Specifically, we compute the Pearson correlation between the estimated and ground-truth rewards in these settings, using IQL+SAC (Garg et al., 2021) as the comparison baseline. The results are presented in Table 7.
In Figure 15, we observe that the variance of the estimated rewards is higher when the ground-truth reward is high. One possible explanation for this high variance in the estimated expert rewards is as follows:
There are multiple equivalent reward formulations that result in optimal trajectories, and the maximum entropy objective selects the one with the highest entropy. Our actor-critic architecture, optimized with the maximum entropy inverse RL objective, leads to a more evenly distributed reward structure for expert demonstrations. Consequently, rewards closer to the expert tend to exhibit higher variance, a phenomenon also observed in (Freund et al., 2023).
Figure 15: Reward Recovery. The IQ-MPC model successfully recovers rewards in the inverse RL setting, showing a positive correlation with ground-truth rewards. This experiment is conducted on the Cheetah Run task with state-based observations from DMControl.
Method IQL+SAC IQ-MPC (Ours)
Hopper Hop 0.49 0.88
Cheetah Run 0.79 0.87
Walker Run 0.65 0.91
Quadruped Run 0.88 0.93
Table 7: Pearson Correlations of Reward Recovery We evaluate the Pearson correlation between the decoded rewards from IQL+SAC and IQ-MPC in the Hopper Hop, Cheetah Run, Quadruped Run, and Walker Run tasks. Our results demonstrate that IQ-MPC achieves a higher correlation with ground-truth rewards when trained on these tasks.
Appendix F Training Time Evaluation
We evaluate the computational overhead of our approach in comparison to model-free baselines (IQL+SAC (Garg et al., 2021), CFIL+SAC (Freund et al., 2023), HyPE (Ren et al., 2024)) and the model-based baseline (HyPER (Ren et al., 2024)). Additionally, we assess the training time of IQ-MPC with model predictive control enabled and when interacting solely with the policy prior. The experiments are conducted on the Humanoid Walk task, with training time reported in seconds. All baselines are trained using a single RTX 2080 Ti GPU. The results are shown in Figure 16.
Figure 16: Computational Overhead We evaluate the computational cost during training of our model on the Humanoid task. Leveraging a policy prior for direct interaction, instead of relying on MPC, accelerates the training process but may introduce greater instability. Our model requires less computational time compared to HyPER, although its training remains slower than model-free baselines.
Appendix G Additional Theoretical Supports
@@
G.1 Objective Equivalence
In this section, we will provide proof for the reformulation of the second term in Eq.12 for completeness. We borrow the proof from Garg et al. (2021) and slightly modify it to fit our setting with latent representations instead of actual states. The proof is demonstrated in Lemma 2. In Eq.12, we use the mean over encoded latent representation batch sampled from the expert buffer $\mathcal{B}_{E}$ to approximate the mean over initial distribution $\tilde{p}_{0}$ on latent representation.
Lemma 2 (Objective Equivalence).
@@ G.1 Objective Equivalence
G.2 Policy Update Guarantee
@@
G.3 Analysis on the Consistency Loss
We provide a more detailed analysis regarding the relationship between minimizing the consistency loss in Eq.11 and minimizing T2 in the bound provided by Lemma 1.Specifically, our consistency loss directly minimizes the upper bound of T2 under following assumptions:
Assumption 1.
The latent dynamics $d:\mathcal{Z}\times\mathcal{A}\rightarrow\Delta_{\mathcal{Z}}$ is approximately a Gaussian distribution on latent space $\mathcal{Z}$ with $\mathcal{N}(\mu_{d},\sigma_{d}^{2})$.
Assumption 2.
The standard deviation of our learned latent dynamics $\hat{\sigma}_{d}$ is close to the actual standard deviation $\sigma_{d}$.
Considering T2 in Lemma 1 and neglecting the constant coefficient, according to Pinsker Inequality, we have:
$\mathbb{E}_{\rho^{\pi}_{\hat{\mathcal{M}}}}\Big{[}D_{TV}(d(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}),\hat{d}(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}))\Big{]}\leq\mathbb{E}_{\rho^{\pi}_{\hat{\mathcal{M}}}}\sqrt{\frac{1}{2}D_{KL}(d(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}),\hat{d}(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}))}$
With Assumption 1 and 2, we can represent the KL divergence by mean and standard deviation of actual and learned latent dynamics:
$D_{KL}(d(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}),\hat{d}(\mathbf{z}^{\prime}|\mathbf{z},\mathbf{a}))=\log\frac{\hat{\sigma}_{d}}{\sigma_{d}}+\frac{\sigma_{d}^{2}+(\mu_{d}-\hat{\mu}_{d})^{2}}{2\hat{\sigma}_{d}^{2}}-\frac{1}{2}\approx\frac{(\mu_{d}-\hat{\mu}_{d})^{2}}{2\sigma_{d}^{2}}$
Given a predicted latent state $\hat{z}^{\prime}$ from the learned dynamics $\hat{d}$ and an actual latent state $z^{\prime}=h(s^{\prime})$ encoded from the next state observation with unknown dynamics $d$, minimizing the L2 loss approximately minimizes the distance between the means of the learned and actual latent dynamics distributions. This, in turn, minimizes the right-hand side of the Pinsker inequality under our assumptions. Consequently, our consistency loss minimizes the statistical distance between the dynamics.